OpenStudy (anonymous):

What is the equation of the line that passes through the points (0, 5) and (2, 3)?

OpenStudy (anonymous):

HELPP!

OpenStudy (texaschic101):

first we find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (0,5)...x1 = 0 and y1 = 5 (2,3)... x2 = 2 and y2 = 3 now we sub slope(m) = (3 - 5) / (2 - 0) slope(m) = -2/2 slope(m) = -1 now we will use y = mx + b and since we have found the slope(m), we need to find the y intercept (b). y = mx + b slope(m) = -1 you can use either of your points.. (0,5)...x = 0 and y = 5 now we sub 5 = -1(0) + b 5 = b so your equation is : y = -1x + 5 or just y = -x + 5 any questions ?

OpenStudy (anonymous):

NOO THANKS